OpenStudy (anonymous):

pleeeeeeeeeeeeaaaaaaaase helllllp me.......... a certain machine make matches. one match in 10000 on average is defective. using a suitable approximation, find the probability that a random sample of 45000 matches will include 2, 3 or 4 defective matches

OpenStudy (anonymous):

@nincompoop

OpenStudy (anonymous):

@IrishBoy123

OpenStudy (anonymous):

@wio

OpenStudy (anonymous):

yesss.. actually am learning poisson distribution

OpenStudy (anonymous):

and ?

OpenStudy (alekos):

P(2 defects) = P(1 defect) x P(1 defect) P(3 defects) = P(2 defects) x P(1 defect) P(4 defects) = P(3 defects) x P( 1 defect)

OpenStudy (anonymous):

how would you calculate them ?

OpenStudy (alekos):

P(2 or 3 or 4 defects) = P(2 defects) + P(3 defects) + P(4 defects)

OpenStudy (anonymous):

help me with the numbers

OpenStudy (alekos):

well you already know that P(1 defect) = 1/10000 so from there it is just simple arithmetic

OpenStudy (anonymous):

yess

OpenStudy (alekos):

use your calculator in exponential or scientific mode

OpenStudy (anonymous):

are you sure its simple arithmetic ? hve u calculated ?

OpenStudy (alekos):

yes, I sure have

OpenStudy (anonymous):

what ans uve got ?

OpenStudy (alekos):

Approximately \[\frac{ 1 }{ 10^{8} }\]

OpenStudy (anonymous):

its wrong.. the answer should be 0.471 its not simple arithmetic

OpenStudy (anonymous):

aniway thanks

OpenStudy (alekos):

can you send me a photo of the actual question?

OpenStudy (anonymous):

it is the actual question

OpenStudy (alekos):

That figure is not possible based on the information you have given

OpenStudy (irishboy123):

i can help you but not whilst others are as that is counter productive for all concerned tag me later

OpenStudy (anonymous):

do you know poisson distribution ?

OpenStudy (alekos):

go for it IrishBoy

OpenStudy (irishboy123):

yes and the answer is 0.471004095.......

OpenStudy (irishboy123):

thx @alekos :p

OpenStudy (irishboy123):

start with the formula and work out \( \lambda \). have you done that? expected outcome given failure rate of 1/10000

OpenStudy (anonymous):

yes

OpenStudy (irishboy123):

then just plug in as per @alekos suggestion ie P(2 or 3 or 4 defects) = P(2 defects) + P(3 defects) + P(4 defects) but what is your \( \lambda\)? that's all that's missing

OpenStudy (irishboy123):

\( \lambda\) is how many you would expect to be defective in a lot of 45,000 given failure rate 1/10,000 then \( P(X=2) = \frac{e^{- \lambda } \lambda^2}{2!}\) do same for 3 and 4 and add and you're done

OpenStudy (anonymous):

thank you

OpenStudy (irishboy123):

i have to go but here are the numbers you should get 2 0.11247859 3 0.168717885 4 0.189807621 \(\Sigma\) 0.471004095

OpenStudy (anonymous):

yeah.. got all